# Uncovering the Functional Roles of Nonlinearity in Memory

### v1

@@ A.1.4 Training on Spike Data
To train on the neural spike data in Sect. 4.5 and Appx. A.3, we embedded the AL-RNN within the Multimodal Teacher Forcing (MTF) framework [10], which enables training directly on non-Gaussian observations through likelihood-based optimization. Specifically, we employed a 1D temporal convolutional encoder to map observed spike trains to the latent space, followed by a Poisson decoder defined as
$\log p(\bm{c}_{t}\mid\mathbf{z}_{t})=\sum_{i}\left(c_{ti}\log\lambda_{ti}-\lambda_{ti}-\log(c_{ti}!)\right)\quad\text{with}\quad\lambda_{t}=\exp(\bm{B}\bm{z}_{t}+\bm{b}),$
where $\bm{z}_{t}$ is the latent state of the AL-RNN and $\lambda_{t}$ the predicted spike rate and $\bm{c}_{t}$ the observed counts. The MTF framework incorporates a latent consistency loss that penalizes the distance between inferred and predicted latent trajectories by
@@
To account for drift in trial-specific spike statistics while preserving a shared task representation, we hierarchically parameterized the decoder following the approach in [11], but applied this structure exclusively to the decoder parameters. Specifically, each trial was associated with a five-dimensional feature vector $\bm{l}^{(j)}$ (with its dimension determined by PCA, Fig.20) that generated trial-specific decoder weights and biases via learned linear projections. The resulting hierarchical Poisson decoder was defined as
@@
This design choice ensures that the AL-RNN learns a consistent dynamic structure across trials. Hierarchizing the RNN itself would allow it to absorb task variation into the dynamics. By restricting flexibility to the output mapping, the AL-RNN remains responsible for solving the core disambiguation problem imposed by the context-dependent task, while the decoder accounts for slow drift or baseline shifts in firing statistics. Although the decoder is not tasked with solving the decision problem, the learned feature vectors nonetheless capture systematic variation linked to task conditions. Specifically, PCA of the trial-specific feature vectors revealed that the first three components were sufficient to linearly separate go vs. no-go trials with $87\%$ accuracy (Fig. 20). A similar level of separability ($\sim 85\%$) was observed when applying PCA directly to the trial-averaged spike data, indicating that the central moments of the firing statistics already encode task-relevant differences. The hierarchical decoder leverages this structure to improve spike reconstruction, but without access to the input structure required to disambiguate the task itself.
@@ A.2 Analysis of Trained Models
To quantify the distribution of latent states across the linear subregions, we extract the last $P$ components of each latent state trajectory and represent their activation patterns as bitcodes, where each component is assigned a binary value based on its sign (1 if positive, 0 otherwise). Formally, the bitcode for a latent state vector $\bm{z}_{t}\in\mathbb{R}^{P}$ at time step $t$ is defined as:
$b_{t}=\sum_{i=1}^{P}\mathbb{I}[z_{t,i}>0]\cdot 2^{P-i},$ (3)
@@
Figure 9: Illustration of the AL-RNN and bitcode assignment. The example displays a 6-dimensional AL-RNN with four linear and two PWL units (left). The PWL units partition the state space into four linear subregions (right). Bitcodes encode positive $(1)$ and negative $(0)$ activation values for these units.
@@
where $\bm{M}(\text{bitcode})$ applies masking to the connectivity matrix $W$, effectively zeroing out the connections of inactive nonlinear units.
@@
using a standard linear solver (implemented with numpy.linalg.solve). If the matrix is singular, no fixed point is considered valid.
To assess the stability of each fixed point, we compute the eigenvalues of the Jacobian matrix:
@@
The maximum Lyapunov exponent, $\lambda_{\max}$, characterizes the average rate of divergence of infinitesimally close trajectories in a dynamical system. It is formally defined as:
@@
[…]
@@ A.3 Datasets
The IMDb dataset [39] consists of 50,000 movie reviews, evenly split between training and test sets, with sentiment labels (positive or negative) for binary classification. It is available on Kaggle at https://www.kaggle.com/datasets/lakshmi25npathi/imdb-dataset-of-50k-movie-reviews. Each review was tokenized and converted to word vectors using pre-trained GloVe embeddings [51]. The sequences are truncated or padded to a fixed length of 128 tokens during training and testing. We did not apply any further text preprocessing beyond tokenization, as the GloVe representations handle most standard normalization. Training and evaluation follow the standard 25,000/25,000 train/test split.
@@
The sequential MNIST (sMNIST) dataset transforms the standard MNIST [34] digit images into sequences by flattening each $28\times 28$ image into a 1D sequence of length 784. We used the dataset as provided in torchvision.datasets [41]. The pixel intensity values are normalized to the range $[0,1]$ before being passed to the AL-RNN. The dataset is split into $60,000$ training sequences and $10,000$ test sequences, following the standard MNIST partitioning.
@@
The Speech Commands dataset [68] consists of 1-second audio recordings with a sampling rate of 16kHz of spoken words from a fixed vocabulary of commands: ["down", "go", "left", "no", "off", "on", "right", "stop", "up", "yes"]. We used the version of the dataset provided by TensorFlow Datasets [1], which includes standardized preprocessing, splitting, and metadata.
@@
The Copy Task is a synthetic memory benchmark where the model is required to store and reproduce a random sequence of discrete symbols after a delay period. Each input sequence consists of a random sequence of symbols represented in one-hot encoding, followed by a blank segment of the same length, and finally a cue signal indicating when to begin recall. The target output is an exact reproduction of the original symbol sequence during the recall phase. Sequence lengths are fixed during training. The training set comprised $1000$ sequences, while the test set contained $200$ sequences. Given that the task is defined over $4$ symbols and $8$ time steps, the total number of unique sequences is $5364^{8}=65,$ 536. This represents a substantial combinatorial space, ensuring that the relatively small training set could not fully cover the distribution of possible sequences, thereby preventing the model from simply memorizing the training examples. Reported test accuracies reflect the percentage of correctly predicted symbols in the test sequences after the delay period.
@@
In the Addition Problem, popularized in [28], the goal is to sum the values at the marked positions and output the result at the end of the sequence. This task tests the model’s ability to integrate sparse, context-dependent information across long temporal spans. The two randomly chosen indices indicating the values to be summed are presented in the first half of the sequence. The AL-RNN is trained to produce the sum of the two digits at the final time step. The training set comprised $2000$ examples, while the test set contained $200$ examples. Reported test errors quantify the mean-squared error between the predicted and the correct sum.
@@
The Contextual Multistability task is a custom-designed integration task inspired by neuroscience paradigms of context-dependent decision-making [42]. Each trial begins with a one-hot context cue, indicating which integration policy the model should adopt. A sequence of noisy evidence is presented across several time steps, and the model is required to integrate this evidence according to the initial context. At the final time step, a recall cue triggers the model to produce a decision based on the integrated evidence. If the initial context is inverted, the decision boundary is reversed. The AL-RNN is trained to predict the correct label based on its final latent state. We trained on $1000$ example sequences, and tested on $200$ sequences.
@@
The CRCNS PFC-1 dataset [55], includes single-unit recordings from medial prefrontal cortex and auditory cortex of rats performing a context-dependent stimulus selection task. We focused on the longest available session (Day 4) from the first rat (CR12B), comprising approximately 900 trials. For this rat, only auditory cortex (AC) neurons were recorded. Each trial involves auditory stimulus comprising a pitch component (high or low warble, presented bilaterally) and a spatial cue (broadband noise from left or right). The correct response—"go" or "no-go"—depends on the current task rule: either localization (respond based on side) or pitch discrimination (respond based on pitch). The task is structured in repeating blocks. The task was structured into repeating blocks of four trial types that defined the task stage.
@@
The SCAN dataset [33] is a synthetic benchmark designed to evaluate systematic compositional generalization in sequence-to-sequence learning. In SCAN, natural language commands are translated into sequences of low-level actions in a navigation environment. For example, the command "jump twice" is mapped to the action sequence "JUMP JUMP", while a more complex command like "walk around right" is interpreted as "RTURN WALK RTURN WALK RTURN WALK RTURN WALK". The mapping is deterministic and unambiguous, with each command generated by a phrase-structure grammar that defines valid language-action mappings. The commands are constructed from a small set of primitives:
@@ A.4 Further Results
A.4.1 Speech Commands
To verify that the trends observed in sMNIST generalize across modalities, we evaluated the AL-RNN on the Speech Commands dataset [68], a ten-class audio classification task using spoken commands sampled at $16$ kHz.
We used the same evaluation setup as for sMNIST, with classification based on the final latent state after processing the input sequence. As in the visual domain, we observed that increasing nonlinearity improved classification accuracy up to an intermediate number of nonlinear units, after which performance declined for fully nonlinear models (Fig. 11a). Latent states from models with moderate nonlinearity again exhibited clear class-specific clustering in PC space (Fig. 11b), and bitcode alignment within classes remained high (Fig. 13)—closely mirroring the gating behavior observed in sMNIST (Sect. 4.1). These results confirm that the latent integration and partitioning mechanisms learned by the AL-RNN generalize robustly across visual and auditory modalities.

### v2

@@ A.1.4 Training on Spike Data
To train on the neural spike data in Sect. A.4.5 and Appx. A.3, we embedded the AL-RNN within the Multimodal Teacher Forcing (MTF) framework (Brenner et al., 2024b), which enables training directly on non-Gaussian observations through likelihood-based optimization. Specifically, we employed a 1D temporal convolutional encoder to map observed spike trains to the latent space, followed by a Poisson decoder defined as
$\log p(\bm{c}_{t}\mid\mathbf{z}_{t})=\sum_{i}\left(c_{ti}\log\lambda_{ti}-\lambda_{ti}-\log(c_{ti}!)\right)\quad\text{with}\quad\bm{\lambda}_{t}=\exp(\bm{B}\bm{z}_{t}+\bm{b}),$
where $\bm{z}_{t}$ is the latent state of the AL-RNN, $\lambda_{ti}$ the predicted spike rates, and $\bm{c}_{t}$ the observed counts. The MTF framework incorporates a latent consistency loss that penalizes the distance between inferred and predicted latent trajectories by
@@
To account for drift in trial-specific spike statistics while preserving a shared task representation, we hierarchically parameterized the decoder following the approach in (Brenner et al., 2025), but applied this structure exclusively to the decoder parameters. Specifically, each trial was associated with a five-dimensional feature vector $\bm{l}^{(j)}$ (with its dimension determined by PCA, Fig. 23) that generated trial-specific decoder weights and biases via learned linear projections. The resulting hierarchical Poisson decoder was defined as
@@
This design choice ensures that the AL-RNN learns a consistent dynamic structure across trials. Hierarchizing the RNN itself could allow it to absorb task variation into the dynamics. By restricting flexibility to the output mapping, the AL-RNN remains responsible for solving the core disambiguation problem imposed by the context-dependent task, while the decoder accounts for slow drift or baseline shifts in firing statistics.
@@ A.2 Analysis of Trained Models
To quantify the distribution of latent states across the linear subregions, we extract a subset of the PWL units from the full latent state vector. Let $\tilde{\bm{z}}_{t}\in\mathbb{R}^{P}$ denote the last $P$ components of the full latent state at time step $t$, corresponding to the PWL units. We represent their activation patterns as bitcodes, where each component is assigned a binary value based on its sign (1 if positive, 0 otherwise). Formally, the bitcode for $\tilde{\bm{z}}_{t}$ is defined as:
$b_{t}=\sum_{i=1}^{P}\mathbb{I}[\tilde{z}_{t,i}>0]\cdot 2^{P-i},$ (4)
@@
where $\bm{M}(\text{bitcode})$ applies masking to the connectivity matrix $W$, effectively zeroing out the outgoing connections of inactive nonlinear units.
@@
using a standard linear solver (implemented with numpy.linalg.solve). If the matrix is singular, no fixed point is considered valid. To assess the stability of each fixed point, we compute the eigenvalues of the Jacobian matrix:
@@
The maximum Lyapunov exponent, $\lambda_{\max}$, characterizes the average rate of divergence of infinitesimally close trajectories in a DS. It is formally defined as:
@@
where $\bm{J}_{T-r}$ represents the Jacobian of the system at time step TrT-r, $\left\|\right\|$ denotes the spectral norm, and the product of Jacobians accumulates the local expansion or contraction at each step. For the AL-RNN, these Jacobians are analytically tractable (Eq. 8) within each subregion. To approximate this exponent numerically, we evolved the trained model forward for $5,000$ time steps from a randomly sampled initial condition, discarding initial transients of $500$ time steps. Since the product of Jacobians grows exponentially in chaotic systems (Mikhaeil et al., 2022), we applied the algorithm outlined in (Vogt et al., 2022), which maintains numerical stability by re-orthogonalizing the Jacobian products at regular intervals via QR decomposition.
@@ A.3 Datasets
The IMDb dataset (Maas et al., 2011) consists of 50,000 movie reviews, evenly split between training and test sets, with sentiment labels (positive or negative) for binary classification. It is available on Kaggle at https://www.kaggle.com/datasets/lakshmi25npathi/imdb-dataset-of-50k-movie-reviews. Each review was tokenized and converted to word vectors using pre-trained GloVe embeddings (Pennington et al., 2014). The sequences are truncated or padded to a fixed length of 128 tokens during training and testing. We did not apply any further text preprocessing beyond tokenization, as the GloVe representations handle most standard normalization. Training and evaluation follow the standard 25,000/25,000 train/test split.
@@
The sequential MNIST (sMNIST) dataset transforms the standard MNIST (Lecun et al., 1998) digit images into sequences by flattening each $28\times 28$ image into a 1D sequence of length 784. We used the dataset as provided in torchvision.datasets (maintainers and contributors, 2016). The pixel intensity values are normalized to the range $[0,1]$ before being passed to the AL-RNN. The dataset is split into $60,000$ training sequences and $10,000$ test sequences, following the standard MNIST partitioning.
@@
The Speech Commands dataset (Warden, 2018) consists of 1-second audio recordings with a sampling rate of 16kHz of spoken words from a fixed vocabulary of commands: ["down", "go", "left", "no", "off", "on", "right", "stop", "up", "yes"]. We used the version of the dataset provided by TensorFlow Datasets (71), which includes standardized preprocessing, splitting, and metadata.
@@
The Copy Task is a synthetic memory benchmark where the model is required to store and reproduce a random sequence of discrete symbols after a delay period. Each input sequence consists of a random sequence of symbols represented in one-hot encoding, followed by a blank segment of the same length, and finally a cue signal indicating when to begin recall. The target output is an exact reproduction of the original symbol sequence during the recall phase. Sequence lengths are fixed during training. The training set comprised $1000$ sequences, while the test set contained $200$ sequences. Given that the task is defined over $4$ symbols and $8$ time steps, the total number of unique sequences is $4^{8}=65,536$. This represents a substantial combinatorial space, ensuring that the relatively small training set could not fully cover the distribution of possible sequences, thereby preventing the model from simply memorizing the training examples. Reported test accuracies reflect the percentage of correctly predicted symbols in the test sequences after the delay period.
@@
In the Addition Problem, popularized in (Hochreiter and Schmidhuber, 1997), the goal is to sum the values at the marked positions and output the result at the end of the sequence. This task tests the model’s ability to integrate sparse, context-dependent information across long temporal spans. The two randomly chosen indices indicating the values to be summed are presented in the first half of the sequence. The AL-RNN is trained to produce the sum of the two digits at the final time step. The training set comprised $2000$ examples, while the test set contained $200$ examples. Reported test errors quantify the mean-squared error between the predicted and the correct sum.
@@
The Contextual Multistability task is a custom-designed integration task inspired by neuroscience paradigms of context-dependent decision-making (Mante et al., 2013). Each trial begins with a one-hot context cue, indicating which integration policy the model should adopt. A sequence of noisy evidence is presented across several time steps, and the model is required to integrate this evidence according to the initial context. At the final time step, a recall cue triggers the model to produce a decision based on the integrated evidence. If the initial context is inverted, the decision boundary is reversed. The AL-RNN is trained to predict the correct label based on its final latent state. We trained on $1000$ example sequences, and tested on $200$ sequences.
@@
The CRCNS PFC-1 dataset (Rodgers and DeWeese, 2014), includes single-unit recordings from medial prefrontal cortex and auditory cortex of rats performing a context-dependent stimulus selection task. We focused on the longest available session (Day 4) from the first rat (CR12B), comprising approximately 900 trials. For this rat, only auditory cortex (AC) neurons were recorded. Each trial involves auditory stimulus comprising a pitch component (high or low warble, presented bilaterally) and a spatial cue (broadband noise from left or right). The correct response—"go" or "no-go"—depends on the current task rule: either localization (respond based on side) or pitch discrimination (respond based on pitch). The task is structured in repeating blocks. The task was structured into repeating blocks of four trial types that defined the task stage.
@@
The SCAN dataset (Lake and Baroni, 2018) is a synthetic benchmark designed to evaluate systematic compositional generalization in sequence-to-sequence learning. In SCAN, natural language commands are translated into sequences of low-level actions in a navigation environment. For example, the command "jump twice" is mapped to the action sequence "JUMP JUMP", while a more complex command like "walk around right" is interpreted as "RTURN WALK RTURN WALK RTURN WALK RTURN WALK". The mapping is deterministic and unambiguous, with each command generated by a phrase-structure grammar that defines valid language-action mappings. The commands are constructed from a small set of primitives:
@@
Multi-Task Training
We jointly trained AL-RNNs on 11 cognitive benchmark tasks adapted from Driscoll et al. (2024) presented in randomly interleaved trials. The task suite includes pro- and anti-response variants of delayed response (remember a stimulus direction across a delay, then respond in the same or opposite direction), reaction time (immediately respond in the same or opposite direction as a presented stimulus), and category decision (classify stimulus angle as above or below threshold, then respond with the corresponding or reversed category mapping). We also included match and non-match versions of delayed match-to-sample, where the network must compare two stimuli separated by delays and respond based on whether they match within a 45° tolerance. Two context integration tasks require selective attention to one of two simultaneously presented sensory modalities while integrating noisy evidence across sequential presentations. Finally, a go/nogo task tests simple threshold detection, requiring a response when stimulus amplitude exceeds 0.5 and fixation maintenance otherwise. Together, these tasks probe working memory, response inhibition, selective attention, and context-dependent computation while sharing common temporal structure and input encodings that encourage discovery of reusable computational strategies.
@@ A.4 Further Results
Table 2: Comparison of different models and activations across tasks.
Model / Activation Addition MSE Copy Acc. sMNIST Acc. SpeechCmd Acc.
AL-RNN (GELU) $2\pm 1\cdot 10^{-4}$ $96\pm 3\%$ $8.2\pm 0.2\% 99$$9.2\pm 0.2\%$
AL-RNN (ReLU) $3\pm 1.5\cdot 10^{-4}$ $92\pm 3\%$ $8.2\pm 0.2\% 99$$9.1\pm 0.2\%$
AL-RNN (tanh) $9\pm 3\cdot 10^{-4}$ $90\pm 3\%$ $8.0\pm 0.2\% 99$$9.0\pm 0.2\%$
AL-RNN (hardtanh) $4\pm 5\cdot 10^{-3}$ $88\pm 6\%$ $7.9\pm 0.3\% 99$$9.0\pm 0.1\%$
LSTM $8.8\pm 10\cdot 10^{-4}$ $54\pm 12\%$ $8.5\pm 0.2\% 99$$9.1\pm 0.1\%$
GRU $6.3\pm 10\cdot 10^{-4}$ $69\pm 7\%$ $8.0\pm 0.2\% 99$$9.0\pm 0.1\%$
Table 3: Performance across tasks for different activation types and sparsity levels.
Model Copy Acc. Addition MSE Contextual Integration Acc.
Linear ($P=0/50$) $58\pm 3\%$ $0.15\pm 0.01$ $\approx 50\%$
Sparse (GeLU, $P=3/50$) $96\pm 3\%$ $2\pm 1\cdot 10^{-4}$ $\approx 95\%$
Sparse (ReLU, $P=3/50$) $92\pm 3\%$ $3\pm 1.5\cdot 10^{-4}$ $\approx 95\%$
Sparse (tanh, $P=3/50$) $90\pm 3\%$ $9\pm 3\cdot 10^{-4}$ $\approx 95\%$
Sparse (hardtanh, $P=3/50$) $88\pm 3\%$ $4\pm 5\cdot 10^{-3}$ $\approx 95\%$
Full (GeLU, $P=50/50$) $38\pm 6\%$ $6\pm 2\cdot 10^{-4}$ $\approx 95\%$
Full (ReLU, $P=50/50$) $39\pm 5\%$ $7\pm 3\cdot 10^{-4}$ $\approx 95\%$
Full (tanh, $P=50/50$) $58\pm 6\%$ $3.6\pm 2\cdot 10^{-3}$ $\approx 95\%$
Full (hardtanh, $P=50/50$) $46\pm 4\%$ $6\pm 5\cdot 10^{-3}$ $\approx 95\%$
A.4.1 Impossibility of Solving the Addition Problem with Linear RNNs
Proposition 1.
An AL-RNN without ReLU nonlinearity cannot solve the addition problem.
Proof.
Consider a linear RNN where $\phi(z)=z$, yielding the dynamics
$z_{t}=\tilde{A}z_{t-1}+Cs_{t}+h,$ (10)
where $\tilde{A}=A+W$ and $s_{t}=(s_{1,t},s_{2,t})^{\top}$. The output at time $T$ is given by
$y_{T}=Bz_{T}=B\left(\tilde{A}^{T}z_{0}+\sum_{t=1}^{T}\tilde{A}^{T-t}(Cs_{t}+h)\right).$ (11)
This can be rewritten as a linear functional of all inputs:
$y_{T}=\sum_{t=1}^{T}\left(\alpha_{t}s_{1,t}+\beta_{t}s_{2,t}\right)+\gamma,$ (12)
where $\alpha_{t}=B\tilde{A}^{T-t}C_{:,1}$, $\beta_{t}=B\tilde{A}^{T-t}C_{:,2}$, and $\gamma$ includes contributions from $z_{0}$ and $h$. Crucially, the coefficients $\alpha_{t}$ and $\beta_{t}$ are fixed functions of time, independent of the input data. For the addition problem, the target output must satisfy $y_{T}=s_{1,t_{1}}+s_{1,t_{2}}$, where $t_{1}$ and $t_{2}$ are the (random) time points at which $s_{2,t_{1}}=s_{2,t_{2}}=1$. This requires the effective coefficient of $s_{1,t}$ in the output to be
$\alpha_{t}^{\text{eff}}=\begin{cases}1&\text{if }s_{2,t}=1\\ 0&\text{if }s_{2,t}=0\end{cases}.$ (13)
However, in the linear system, $\alpha_{t}$ is fixed and cannot depend on $s_{2,t}$. Since $t_{1}$ and $t_{2}$ are randomly positioned across trials, no fixed sequence $\{\alpha_{t}\}_{t=1}^{T}$ can satisfy $\alpha_{t_{1}}=\alpha_{t_{2}}=1$ and $\alpha_{t}=0$ for $t\notin\{t_{1},t_{2}\}$ for all trials simultaneously. Therefore, a linear PLRNN cannot implement the conditional integration required by the addition problem, as it lacks the ability to gate the contribution of $s_{1,t}$ based on the value of $s_{2,t}$.
A.4.2 Classification Tasks
sMNIST
In the sMNIST task, $28\times 28$ digit images are flattened into a sequential time series of length 784, with one pixel input per time step. Classification is based solely on the final latent state. Bitcode analysis provides a clear mechanism by which nonlinearity increases performance: each class aligns with its own set of closely neighboring bitcodes. For instance, for $P=50$, the average variation in the bitcode within each class was just $0.06\pm 0.02$ (see Fig. 10, 0.5 indicates chance level), meaning that on average $48$ out of $50$ PWL units had the same sign at the final state. Between classes, PWL units use different bitcodes. We observed that switching between these codes during sequential processing is aligned with the onset of visually informative stroke patterns (see Fig. 11 for example illustrations). This allows the AL-RNN to implement class-specific linear integration dynamics: while all digit classes are processed through slow modes with maximum eigenvalues close to 1 (Fig. 9c), subtle class-dependent differences in these eigenvalues, arising from the distinct linear subregions occupied by each class, implement differentiated integration rates. Since the input projection itself is linear (via the C matrix), it is the autonomous dynamics within each subregion that reshape the latent manifold: local flows with slightly divergent eigenvalues drive trajectories for different digit classes apart in latent space, resulting in well-separated clusters that facilitate discrimination without requiring the input integration mechanism itself to be nonlinear (Figs. 2f &\& 12).
Figure 9: a: Mean latent trajectories of the first PC for each digit class in the sMNIST task. The trajectories reveal a slow accumulation of evidence along the primary axis, similar to sentiment classification (Fig. 2c). Interestingly, distinct branching patterns are visible: the digit ’6’ diverges first, reflecting its characteristic structure in the initial image segments. Digits ’8’ and ’9’ share overlapping paths until around the midway point, aligning with their visual similarity in the upper halves of the images. b: Explained variance ratio of the first $15$ PCs indicates a more complex structure compared to the $2D$ sentiment attractor in Fig. 2. (c) Maximum eigenvalue analysis across digit classes for models with varying nonlinearity. Linear and minimally nonlinear models ($P=1,5$) maintain maximum eigenvalues consistently near $1.0$ across all digits, indicating that a uniform slow mode is shared across all classes. Models with moderate nonlinearity ($P=50$) preserve eigenvalues close to unity but show class-specific variations, demonstrating that nonlinearity enables subtle, digit-specific modulation of the slow manifold while maintaining the overall slow integration character. Fully nonlinear models ($P=100$) exhibit systematically larger eigenvalues across all classes, providing a mechanistic explanation for the overdispersion and reduced performance observed at high $P$.
Figure 11: Analysis of bitcode activations for the sMNIST task. Each panel shows the average Hamming distance from the reference state $(0,0,0)$ for a model with $P=3$ bitcode activations, aligned across time for all 10 digits. The PWL units switch primarily during visually informative stroke segments, effectively gating nonlinear processing. Digits with sharp vertical onsets (e.g., “1”, “4” and “6”) display early bitcode activations, whereas digits with round strokes (e.g., “0”, “8”) activate later. Digits with similar overall shapes (e.g., “8” vs. “9”) diverge in their bitcode representation toward the end of the sequence. These qualitative differences illustrate how the different linear subregions of the AL-RNN support the dynamic discrimination of the different digits.
Figure 12: Analysis of class manifolds for the sMNIST task. (a) Kernel Density Estimates of key metrics across nonlinear configurations ($P=0,50,100$). The fully linear model ($P=0$) exhibits tightly packed, uniform class variance (low Gini coefficients and Max-Min ratios). Moderate nonlinearity ($P=50$) achieves optimal balance with $CV\approx 1$, producing spatially distinct manifolds. The fully nonlinear model ($P=100$) shows high Gini coefficients and wide Max-Min ratios, indicating uneven variance distribution where some manifolds are over-compressed while others are spread out. (b) Final latent states projected onto the first three PCs for frequently confused digits 3 (orange) and 5 (green). At $P=50$, manifolds are spatially distinct and well-separated. At $P=100$, they become elongated and entangled, with digit 5 particularly over-compressed, directly correlating with observed misclassification patterns.
Speech Commands